OpenStudy (anonymous):

find the derivative of: sqrt(2x) Book says f'(x)= sqrt(2)/(2*sqrt(x)) I can't get that.

OpenStudy (anonymous):

its 2x^.5 so \[(.5)(2)x^(.5-1) = x^(-.5)\]

OpenStudy (anonymous):

then, x^-.5 = \[1/sqrtx\]

OpenStudy (anonymous):

idk y your book says that

OpenStudy (anonymous):

That's what I've been getting, too. I even tried deriving via the differential and couldn't come to the book's answer.

OpenStudy (anonymous):

*differential equation.

OpenStudy (anonymous):

is there another step in the instructions? but thats the derivative fersure

OpenStudy (anonymous):

In Exercise 1 through 28, differentiate the given function. Simplify your answers. \[y=\sqrt{2x}\]

OpenStudy (anonymous):

oh. then idk what to tell you. thats the answer tho. just one problem dude. just move on. YOU know your right.

OpenStudy (phi):

one way to see the answer: \[\sqrt{2x}= \sqrt{2}x^{\frac{1}{2}}\] and the derivative is \[\sqrt{2} \cdot \frac{1}{2}x^{-\frac{1}{2}}\]

OpenStudy (phi):

In other words \[\sqrt{2}\frac{d}{dx}x^{\frac{1}{2}}= \sqrt{2}\frac{1}{2}x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

^^ nice

OpenStudy (phi):

The other way is to use the chain rule

OpenStudy (phi):

\[\frac{d}{dx}(2x)^{\frac{1}{2}}= \frac{1}{2}(2x)^{-\frac{1}{2}}\cdot 2=\frac{1}{\sqrt{2}{\sqrt{x}}}= \frac{\sqrt{2}}{2\sqrt{x}}\]

OpenStudy (anonymous):

I wasn't looking at \[\sqrt{2}\] as a constant. Thank you!